How do you show that (8^n+3^n)/(8^-n+3^-n)=24^n without using the logarith way?
Answer by rothauserc(4718)

(8^n + 3^n) / (8^-n + 3^-n)
work with the denominator
(8^-n + 3^-n) = (1/8^n + 3^n)
LCM is 24, therefore
(8^-n + 3^-n) = (3^n + 8^n) / 24^n
invert denominator and multiply by numerator
(8^n + 3^n) * 24^n / (8^n + 3^n) = 24^n
note that denominator and numerator cancels (8^n + 3^n)